3x(2x-5)(3x+1) = 0
Answer by jmorgan22(8)

3x(2x-5)(3x+1)=0
If any individual factor on the left-hand side of the equation is equal to 0, the entire expression will be equal to 0.
3x=0
(2x-5)=0
(3x+1)=0

Set the first factor equal to 0 and solve.
3x=0

Divide each term in the equation by 3.
3x/3=0/3

Simplify the left-hand side of the equation by canceling the common factors.
x=0/3

0 divided by any number or variable is 0.
x=0

Set the next factor equal to 0 and solve.
(2x-5)=0

Remove the parentheses around the expression 2x-5.
2x-5=0

Since -5 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 5 to both sides.
2x=5

Divide each term in the equation by 2.
2x/2=5/2

Simplify the left-hand side of the equation by canceling the common factors.
x=5/2

Set the next factor equal to 0 and solve.
(3x+1)=0

Remove the parentheses around the expression 3x+1.
3x+1=0

Since 1 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 1 from both sides.
3x=-1

Divide each term in the equation by 3.
3x/3=-1/3

Simplify the left-hand side of the equation by canceling the common factors.
x=-1/3

The final solution is all the values that make 3x(2x-5)(3x+1)=0 true.
Answer: x=0, 5/2, -1/3